How many titanium are in a pure titanium bicycle frame with a mass of 1.28kg
1
Expert's answer
2017-02-17T07:40:46-0500
n = m/M
M (Ti) = 47.9 g/mol
n (Ti) = 1280/47.9 = 26.7 mol